OpenStudy (anonymous):

Solve –3(2x – 1) = x + 10. You must show all work

OpenStudy (anonymous):

I gotchu !

OpenStudy (anonymous):

-3(2x-1)=x+10 -6x-3=x+10 -7x-13=0 -7x=13 x=-13/17

OpenStudy (anonymous):

sorry its -13/7

OpenStudy (anonymous):

thanks